# A minimal coalition logic

### v2

@@ Appendix A Proofs for some results
  • $\mathtt{A}\text{-}\mathtt{MG}$. It suffices to show $\vdash_{1}(\langle[\emptyset]\rangle(\phi\rightarrow\psi)\land\langle[\mathrm{A}]\rangle\phi)\rightarrow\langle[\mathrm{A}]\rangle\psi$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}(\langle[\emptyset]\rangle(\phi\rightarrow\psi)\land\langle[\mathrm{A}]\rangle\phi)\rightarrow\langle[\mathrm{A}]\rangle((\phi\rightarrow\psi)\land\phi)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle[\mathrm{A}]\rangle((\phi\rightarrow\psi)\land\phi)\rightarrow\langle[\mathrm{A}]\rangle(\phi\land\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Mon}$, $\vdash_{1}\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow\langle[\mathrm{A}]\rangle\psi$. Then, $\vdash_{1}(\langle[\emptyset]\rangle(\phi\rightarrow\psi)\land\langle[\mathrm{A}]\rangle\phi)\rightarrow\langle[\mathrm{A}]\rangle\psi$.
@@
  • $\mathtt{A}\text{-}\mathtt{MC}$. Assume $\mathrm{A}\subseteq\mathrm{B}$. We want to show $\vdash_{1}\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\mathrm{B}]\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{1}\langle[\mathrm{B}-\mathrm{A}]\rangle\top$. Then, $\vdash_{1}\langle[\mathrm{A}]\rangle\phi\rightarrow(\langle[\mathrm{A}]\rangle\phi\land\langle[\mathrm{B}-\mathrm{A}]\rangle\top)$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}\langle[\mathrm{A}]\rangle\phi\land\langle[\mathrm{B}-\mathrm{A}]\rangle\top\rightarrow\langle[\mathrm{B}]\rangle(\phi\land\top)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle[\mathrm{B}]\rangle(\phi\land\top)\rightarrow\langle[\mathrm{B}]\rangle\phi$. Then we can get $\vdash_{1}\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\mathrm{B}]\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Det}$. We want to show $\vdash_{1}\langle[\mathrm{A}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{A}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}(\langle[\mathrm{A}]\rangle(\phi\lor\psi)\land\langle[\emptyset]\rangle\neg\psi)\rightarrow\langle[\mathrm{A}]\rangle((\phi\lor\psi)\land\neg\psi)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle[\mathrm{A}]\rangle((\phi\lor\psi)\land\neg\psi)\rightarrow\langle[\mathrm{A}]\rangle(\phi\land\neg\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Mon}$, $\vdash_{1}\langle[\mathrm{A}]\rangle(\phi\land\neg\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi$. Then, $\vdash_{1}(\langle[\mathrm{A}]\rangle(\phi\lor\psi)\land\langle[\emptyset]\rangle\neg\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Max}$, $\vdash_{1}\neg\langle[\mathtt{AG}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\neg\psi$. Then, $\vdash_{1}(\langle[\mathrm{A}]\rangle(\phi\lor\psi)\land\neg\langle[\mathtt{AG}]\rangle\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi$. Then, $\vdash_{1}(\langle[\mathrm{A}]\rangle(\phi\lor\psi)\rightarrow(\neg\langle[\mathtt{AG}]\rangle\psi\rightarrow\langle[\mathrm{A}]\rangle\phi)$. Then, $\vdash_{1}\langle[\mathrm{A}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{A}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$.
@@
  • $\mathtt{R}\text{-}\mathtt{CN}$. Assume $\vdash_{1}\phi$. We want to show $\vdash_{1}\langle[\mathrm{A}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi$. Note $\vdash_{1}\top\leftrightarrow\phi$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle[\emptyset]\rangle\top\rightarrow\langle[\emptyset]\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{1}\langle[\emptyset]\rangle\top$. Then $\vdash_{1}\langle[\emptyset]\rangle\phi$. Then $\vdash_{1}\langle[\mathrm{A}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Mon}$. Note $\vdash_{2}(\phi\land\psi)\rightarrow\phi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow\langle[\emptyset]\rangle((\phi\land\psi)\rightarrow\phi)$. By Axiom $\mathtt{A}\text{-}\mathtt{MG}$, $\vdash_{2}\langle[\emptyset]\rangle((\phi\land\psi)\rightarrow\phi)\rightarrow(\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi)$. Then $\vdash_{2}\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow(\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi)$. Then $\vdash_{2}\langle[\mathrm{A}]\rangle(\phi\land\psi)\rightarrow\langle[\mathrm{A}]\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Max}$. Note $\vdash_{2}\neg\phi\lor\phi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle[\emptyset]\rangle\top\rightarrow\langle[\emptyset]\rangle(\neg\phi\lor\phi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{2}\langle[\emptyset]\rangle\top$. Then $\vdash_{2}\langle[\emptyset]\rangle(\neg\phi\lor\phi)$. Note $\vdash_{2}\langle[\emptyset]\rangle(\neg\phi\lor\phi)\rightarrow(\langle[\emptyset]\rangle\neg\phi\lor\langle[\mathtt{AG}]\rangle\phi)$ is an instance of Axiom $\mathtt{A}\text{-}\mathtt{Det}$. Then $\vdash_{2}\langle[\emptyset]\rangle\neg\phi\lor\langle[\mathtt{AG}]\rangle\phi$. Then, $\vdash_{2}\neg\langle[\emptyset]\rangle\neg\phi\rightarrow\langle[\mathtt{AG}]\rangle\phi$.
@@
  • $\mathtt{R}\text{-}\mathtt{RE}$. Assume $\vdash_{2}\phi\leftrightarrow\psi$. We want to show $\vdash_{2}\langle[\mathrm{A}]\rangle\phi\leftrightarrow\langle[\mathrm{A}]\rangle\psi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\emptyset]\rangle(\phi\rightarrow\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{MG}$, $\vdash_{2}\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\mathrm{A}]\rangle\psi)$. Then, $\vdash_{2}\langle[\mathrm{A}]\rangle\phi\rightarrow(\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\mathrm{A}]\rangle\psi)$. Then $\vdash_{2}\langle[\mathrm{A}]\rangle\phi\rightarrow\langle[\mathrm{A}]\rangle\psi$. Similarly, we can show $\vdash_{2}\langle[\mathrm{A}]\rangle\psi\rightarrow\langle[\mathrm{A}]\rangle\phi$. Then $\vdash_{2}\langle[\mathrm{A}]\rangle\phi\leftrightarrow\langle[\mathrm{A}]\rangle\psi$.

### v3

@@ 3.1 Concurrent game models
Example 1.
$s_{0}$$s_{1}$$s_{2}$$s_{3}$$s_{4}$$\{l_{a},l_{b}\}$$\{m_{a},l_{a}\}$$\{m_{b},l_{b}\}{}\{\}$$\{m_{a},m_{b},l_{a},l_{b}\}$$\mathtt{(w,n)}$$\mathtt{(n,w)}$$\mathtt{(n,n)}$$\mathtt{(w,w)}$$\mathtt{(n,n)}$$\mathtt{(n,n)}$$\mathtt{(n,n)}$$\mathtt{(n,n)}$
3.2 Language and semantics
In $\mathsf{CL}$, the formula $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$ intuitively indicates some unconditionally available joint action of $\mathrm{A}$ ensures $\phi$, and $\hskip-2.5pt\langle\mathrm{A}\rangle\hskip-2.5pt]\phi$ intuitively indicates every unconditionally available joint action of $\mathrm{A}$ enables $\phi$.
[…]
$\mathtt{M},s\Vdash\Box\phi$ $\Leftrightarrow$ for all $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$, $\mathtt{M},t\Vdash\phi$
$\mathtt{M},s\Vdash\Diamond\phi$ $\Leftrightarrow$ there is $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$ such that $\mathtt{M},t\Vdash\phi$
We use $\models_{\mathsf{CL}}\phi$ to indicate that $\phi$ is $\mathsf{CL}$-valid: $\phi$ is true at all pointed concurrent game models. We often drop $\mathsf{CL}$” when the contexts are clear.
3.3 Axiomatization
Definition 5 (An axiomatic system for $\mathsf{CL}$ [13]).
Monotonicity ($\mathtt{A}\text{-}\mathtt{Mon}$): $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$
Liveness ($\mathtt{A}\text{-}\mathtt{Live}$): $\neg\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\bot$
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}\cup\mathrm{B}]\hskip-2.0pt\rangle(\phi\land\psi)$, where $\mathrm{A}\cap\mathrm{B}=\emptyset$
$\mathtt{AG}$-maximality ($\mathtt{A}\text{-}\mathtt{Max}$): $\neg\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\phi\rightarrow\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\phi$
Inference rules:
Replacement of equivalence ($\mathtt{R}\text{-}\mathtt{RE}$): $\dfrac{\phi\leftrightarrow\psi}{\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\leftrightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi}$
We use $\vdash_{\mathsf{CL}}\phi$ to indicate that $\phi$ is derivable in this system. We often drop $\mathsf{CL}$” when the contexts are clear.
Theorem 1 (Soundness and completeness of $\mathsf{CL}$ [13, 7]).
Definition 6 (Another axiomatic system for $\mathsf{CL}$).
Axioms:
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\phi$, where $\mathrm{A}\subseteq\mathrm{B}$
Liveness ($\mathtt{A}\text{-}\mathtt{Live}$): $\neg\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\bot$
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}\cup\mathrm{B}]\hskip-2.0pt\rangle(\phi\land\psi)$, where $\mathrm{A}\cap\mathrm{B}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$
Inference rules:
Conditional necessitation ($\mathtt{R}\text{-}\mathtt{CN}$): $\dfrac{\phi}{\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi}$
@@ 3.3 Axiomatization
  • $\mathtt{A}\text{-}\mathtt{MG}$. It suffices to show $\vdash_{1}(\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\land\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}(\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\land\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle((\phi\rightarrow\psi)\land\phi)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle((\phi\rightarrow\psi)\land\phi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Mon}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$. Then, $\vdash_{1}(\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\land\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$.
@@
  • $\mathtt{A}\text{-}\mathtt{MC}$. Assume $\mathrm{A}\subseteq\mathrm{B}$. We want to show $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{B}-\mathrm{A}]\hskip-2.0pt\rangle\top$. Then, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{B}-\mathrm{A}]\hskip-2.0pt\rangle\top)$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\land\langle\hskip-2.0pt[\mathrm{B}-\mathrm{A}]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle(\phi\land\top)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle(\phi\land\top)\rightarrow\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\phi$. Then we can get $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{B}]\hskip-2.0pt\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Det}$. We want to show $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{1}(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\land\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle((\phi\lor\psi)\land\neg\psi)$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle((\phi\lor\psi)\land\neg\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\neg\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Mon}$, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\neg\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$. Then, $\vdash_{1}(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\land\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Max}$, $\vdash_{1}\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\psi$. Then, $\vdash_{1}(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\land\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$. Then, $\vdash_{1}(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)$. Then, $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\lor\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\psi)$.
@@
  • $\mathtt{R}\text{-}\mathtt{CN}$. Assume $\vdash_{1}\phi$. We want to show $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi$. Note $\vdash_{1}\top\leftrightarrow\phi$. By Rule $\mathtt{R}\text{-}\mathtt{RE}$, $\vdash_{1}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{1}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$. Then $\vdash_{1}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi$. Then $\vdash_{1}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Mon}$. Note $\vdash_{2}(\phi\land\psi)\rightarrow\phi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle((\phi\land\psi)\rightarrow\phi)$. By Axiom $\mathtt{A}\text{-}\mathtt{MG}$, $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle((\phi\land\psi)\rightarrow\phi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)$. Then $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi)$. Then $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle(\phi\land\psi)\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$.
@@
  • $\mathtt{A}\text{-}\mathtt{Max}$. Note $\vdash_{2}\neg\phi\lor\phi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\neg\phi\lor\phi)$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$. Then $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\neg\phi\lor\phi)$. Note $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\neg\phi\lor\phi)\rightarrow(\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\phi)$ is an instance of Axiom $\mathtt{A}\text{-}\mathtt{Det}$. Then $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\phi\lor\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\phi$. Then, $\vdash_{2}\neg\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\neg\phi\rightarrow\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\phi$.
@@
  • $\mathtt{R}\text{-}\mathtt{RE}$. Assume $\vdash_{2}\phi\leftrightarrow\psi$. We want to show $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\leftrightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$. By Rule $\mathtt{R}\text{-}\mathtt{CN}$, $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)$. By Axiom $\mathtt{A}\text{-}\mathtt{MG}$, $\vdash_{2}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle(\phi\rightarrow\psi)\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi)$. Then, $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow(\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi)$. Then $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$. Similarly, we can show $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi\rightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi$. Then $\vdash_{2}\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\phi\leftrightarrow\langle\hskip-2.0pt[\mathrm{A}]\hskip-2.0pt\rangle\psi$.
4 On seriality, independence of agents, and determinism
As mentioned, concurrent game models have the assumptions of seriality, the independence of agents, and determinism. In this section, against the literature, we argue that the three assumptions do not hold generally. In addition, we also discuss the potential impact of dropping them.
4.1 Coalitions do not have available joint actions in all situations
The real world evolves continuously, but many important artificial settings, such as games, have boundaries, including terminal states. For example, a paper-scissors-rock game terminates when a winner is decided. Intuitively, in terminating states of these artificial settings, agents do not have available actions.
One natural thought is to model termination by introducing a special action, that is, “doing nothing”, which does not change anything. Terminal states are those where every agent can only perform the special action. We do not see that this causes a problem, although it seems conceptually incorrect. Note that some literature, such as [2], acknowledges that “doing nothing” may lead to state change.
Dropping seriality might change things. Take Alternating-time Temporal Logic $\mathsf{ATL}$ as an example, whose models are concurrent game models. In $\mathsf{ATL}$, paths, that are infinite sequences of states, play a crucial role. If we drop seriality, we shall consider both finite paths and infinite paths.
4.2 Available joint actions are not unconditional in all situations
Agents’ actions often interfere with each other, and whether a coalition can perform an action is often conditional on other agents’ actions at the same time. The interference among agents’ actions can be due to many reasons, including a lack of resources. It fails the independence of agents. If we want to consider interactions among agents’ actions, we should relax the assumption of agent independence.
What follows are three examples where the independence of agents does not hold; the last two are from Sergot [16].
Example 2.
Example 3.
Example 4.
There are two agents, $a$ and $b$, and a precious vase. The vase can be in one of three locations: $\mathtt{in}$, $\mathtt{out}$, and $\mathtt{over}$. It might be raining or not, which is out of the control of $a$ and $b$. It is wrong for the vase to be $\mathtt{out}$ when it is raining. Agent $a$ can move the vase between $\mathtt{in}$ and $\mathtt{out}$, and agent $b$ can move it between $\mathtt{out}$ and $\mathtt{over}$. The vase is in $\mathtt{out}$, and it is not raining. In this case, agent $a$ can move the vase to $\mathtt{in}$, agent $b$ can move the vase to $\mathtt{over}$, but they cannot do it simultaneously.
One natural thought is to treat the actions in the scenarios given above as trying to do something. Since people can always try to do things at the same time, the independence of agents holds. There are some issues with this idea. First, it would trivialize the notion of availability: all joint actions would be available. However, availability is not a trivial notion. For example, it makes perfect sense to say “they cannot sit at the same time” in the first example. Second, this thought seems incompatible with determinism. For example, it is natural to think that the outcome of $a$ and $b$ trying to sit is not deterministic in the first example. Finally, Royakkers and Hughes [15] found that this idea can lead to implausible conclusions in certain situations regarding whether an agent is accountable for a result.
There are advantages to dropping the assumption of agent independence: it lets us model conditional abilities. Roughly, conditional abilities are something like this: given that the coalition $\mathrm{A}$ is going to behave in a certain way, the coalition $\mathrm{B}$ is able to ensure a certain goal. We refer to [6] for more discussion of conditional abilities.
There are many situations where some joint actions of all behaving agents have more than one outcome state. What follows are some examples; the last two of them are from [16].
Example 5.
Example 6.
Example 7.
Furthermore, in many situations, we consider some but not all behaving agents, who form the grand coalition. The reason could be that we do not know exactly who is behaving, or that we know who is behaving, but only some of them are important to us. In these situations, determinism fails.
A natural idea is to suppose an implicit agent, the environment, which represents nature, unknown agents, and ignored agents as a whole. About this idea, let us quote some words from Sergot [17]: “In many examples however the device of thinking of the environment as a kind of agent is very much less convincing. Suppose that $a$ and $b$ can throw the vase out of a window. If it is thrown out of the window and breaks, the effects of the environment are to contribute—what? fragility? the angle at which the vase strikes the ground? In examples, common in the literature on causation, where agents throw rocks at a bottle with imperfect aim, what is the environment’s contribution? To affect the trajectory of the throw? To affect the trajectory of each throw differently? To affect the trajectory of each throw differently depending on who else throws at the same time? To contribute fragility to the bottle? In all but the very simplest examples the device seems extremely contrived, and does not scale.
Remarks
The granularity of actions can play a role in whether determinism holds. For example, in Example 7, we could specify which ways of trying to pass can let $a$ pass through and which ways cannot. Then, the joint actions of $a$ and $b$ would be deterministic.
5.1 General concurrent game models
Definition 7 (General concurrent game models).
  1. (1)
  2. (2)
  3. (3)